I am trying to figure out how to elimination problems....
Not having much luck....can someone help me, point me in the right direction?
6x+2y-4=0 and 10x+7y=-8

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Here's one I just did, see if you can do it like this.
If you have trouble, or to check your work, email me at
gsihoutx@aol.com
Solve the system by the elimination method.
5x + 5y = -13
7x - 3y = 19
-------------
To eliminate one of the variables, multiply one or both eqns to get the same coefficient for one of them. You can do either, I'll do y
--------
5x + 5y = -13 x3 --> 15x + 15y = -39
7x - 3y = 19 x5 --> 35x - 15y = 95
--------------------------------------
15x + 15y = -39
35x - 15y = 95
--------------- Add
50x = 56
x = 56/50 = 1.12
-----------------
Usually it's "elimination and substution", once you find one, you can sub and find the others.
7x - 3y = 19
7*1.12 - 3y = 19
3y = 7.84 - 19 = - 11.16
y = -3.72

Answer by stanbon(75887)

6x+2y-4=0
10x+7y=-8
-----------
Rearrange
6x + 2y = 4
10x+ 7y = -8
--------------------
Multiply thru the 1st Eq. by 7
Multiply thru the 2nd Eq. by 2
--------------------
42x + 14y = 28
20x + 14y = -16
------------------------
Subtract the 2nd equation from the 1st:
22x = 44
x = 2
--------------------
Subtract x=2 into 6x+2y = 4 to solve for "y":
6*2 + 2y = 4
2y = -8
y = -4
==========================
Cheers,
Stan H.